### v3

# LEMON: A Large Endoscopic MONocular Dataset and Foundation Model for Perception in Surgical Settings

Abstract
Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos and less than 30 hours of footage, which leads to poor model generalization. To address this constraint, a new dataset called LEMON has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage across multiple procedure types, LEMON offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel downstream tasks. To demonstrate the effectiveness of this diverse dataset, we introduce LemonFM, a foundation model pretrained on LEMON using a novel self-supervised augmented knowledge distillation approach. LemonFM consistently outperforms existing surgical foundation models across four downstream tasks and six datasets, achieving significant gains in surgical phase recognition (+9.5pp, +9.4pp, and +8.4pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), surgical action recognition (+4.4pp of mAP in CholecT50), surgical tool presence detection (+5.3pp and +10.2pp of mAP in Cholec80 and GraSP), and surgical semantic segmentation (+8.3pp of mDice in CholecSeg8k). LEMON and LemonFM will serve as foundational resources for the research community and industry, accelerating progress in developing autonomous robotic surgery systems and ultimately contributing to safer and more accessible surgical care worldwide. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/LEMON.
1 Introduction
Figure 1: Performance comparison between LemonFM, our foundation model pretrained on LEMON, and other surgical foundation models. The numeric labels mark the axis ticks. The full results are available in Table 3 and 4.
Figure 2: Data curation pipeline. (a) Surgical videos are collected from YouTube. (b) The videos are summarized as storyboards (trivial to annotate and classify). We train a storyboard classifier to identify videos with rich surgical content, and manually verify the selected videos. (c) During the video selection and trimming phase, a trained frame classifier categorizes each frame as either surgical or non-surgical, allowing us to trim non-surgical frame segments at the beginning and end of each video. We further filter videos by requiring at least 90% of their frames to be surgical frames. (d) In the video preprocessing stage, we obliterate non-surgical frames and non-surgical regions within surgical frames. (e) During the video annotation phase, we utilize the video title as a primary cue to determine the procedure and surgery type. In cases where the video titles do not explicitly include any of our procedure name keywords, we employ ChatGPT to match the titles with surgical procedure types. Both the final videos and their corresponding labels are manually quality controlled.
Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [24]. It promises safer and more consistent operations, fewer complications, shorter operative times, lighter workloads for surgeons, and wider access to high‑quality care. To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38]. However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [54, 40, 44]. The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [17, 2]. The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [47]. These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
Recent advances in self-supervised learning have opened up new research avenues. Pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [46]. Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks. By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical vision analysis.
Table 1: Comparison of LEMON’s size to other existing public surgical datasets.
Dataset Procedure Video Hour Frame Frame (1 fps)
M2CAI16 1 41 262626 2.42.42.4M  [^95] K
Cholec80 1 80 515151 4.64.64.6M 183183183K
CholecT50 1 50 282828 2.52.52.5M 101101101K
AutoLaparo 1 21 232323 2.12.12.1M  [^83] K
GraSP 111 131313 333333 3.53.53.5M 117117117K
GenSurgery 282828 318231823182 680680680  [^70] M 2.42.42.4M
SurgeNetXL 232323 325332533253 576576576  [^52] M 2.12.12.1M
LEMON $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
Figure 3: Comparison between LEMON and other surgical datasets sourced from the web. Video segments from each dataset illustrating the data curation differences (left). Procedures which are not covered by both GenSurgery and SurgeNetXL or any other public dataset (right).
Contributions. To address the scarcity of high‑quality surgical data and the practical difficulty of collecting it, we propose a novel multi-stage data curation pipeline designed specifically for online surgical videos (Fig. 2). In contrast to previous efforts that relied on limited, manually curated validation processes with datasets containing non-surgical content, our pipeline employs a systematic approach to filter out non-surgical videos, remove non-surgical content from surgical videos, and verify the curated surgical videos and their annotations. Building upon this data curation pipeline, we present LEMON, a large-scale dataset comprising more than 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage (Table 1). Additionally, we propose two novel supervised video classification tasks based on the accompanying annotations: multi-class surgical procedure classification and binary surgery type classification, and establish a leaderboard for them. To the best of our knowledge, LEMON is the largest open access surgical dataset to date, both in total frames and videos and in the breadth of distinct surgical procedures it includes. To model the features of surgical frames, we introduce LemonFM, a foundation model pretrained on LEMON that leverages self-supervised learning with our proposed augmented knowledge distillation method. We evaluated LemonFM on six datasets: AutoLaparo [52], Cholec80 [51], M2CAI16 [51], GraSP [4], CholecT50 [43], and CholecSeg8k [23], covering four tasks including phase recognition, tool presence detection, action recognition, and semantic segmentation. Under both linear probing and full fine-tuning protocols, LemonFM consistently surpasses existing surgical foundation models and task-specific architectures (Fig. 1, Table 2, 3, 4). Moreover, LemonFM retains these advantages even when fine‑tuned with only half of the labeled data. These results showcase LemonFM’s strong generalization and data efficiency over diverse surgical scenarios and underscore the value of LEMON as a broadly useful resource for the community. We provide open access to our dataset, models, and code for data curation and training.
2 Related Work
Surgical datasets. Previous works have released open-access datasets aimed at addressing specific surgical vision tasks and procedures (Table 1). Despite valuable, these datasets suffer from a narrow scope of surgical procedures and limited dataset sizes. GenSurgery [48] and SurgeNetXL [25] expand procedural coverage by compiling public videos, but the lack of a curation step to remove the non-surgical content from the public videos (Fig. 3) might introduce spurious features that can confound discriminative models. Additionally, several private datasets exist, such as SVL [57] and SAGES CVS [45], but their restricted access limits their utility and broader adoption. To overcome these limitations, we introduce a novel data curation pipeline that yields LEMON, the largest open-access surgical dataset to date, surpassing previous datasets in both the number of frames and videos and the diversity of distinct surgical procedures represented.
Self-supervised learning. Self-supervised learning (SSL) has gained significant attention due to its ability to reduce reliance on large-scale annotated datasets, a benefit that is particularly relevant for surgical applications. Recently, Ramesh et al. [46] investigated the effectiveness of different SSL approaches for surgical applications. Their experimental results show that DINO [8] outperformed MoCov2 [11], SwAV [7], and SimCLR [10] in surgical applications. As part of a broader effort to advance self-supervised surgical models, Hirsch et al. [22] built a private surgical dataset and trained a Masked Siamese Networks (MSN) [3] on it. Despite encouraging results, the work’s wider applicability and reproducibility are limited by its reliance on private data and a relatively narrow evaluation scope: surgical phase recognition on Cholec80 and polyp characterization on PolypSet [33]. Similarly, in Endo-FM [53], the authors proposed to train a student-teacher foundation model on a composite dataset that combined private and public data sources. Although showing promising results, the approach is also of limited accessibility. More recently, Batic et al. [5] created a dataset by merging nine public datasets. They trained a model based on masked autoencoders (MAE) [21], but showed limited performance, likely due to their reduced dataset size.
3 Proposed LEMON Dataset
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need, we propose LEMON, a dataset comprising more than 4K videos totaling 938 hours of surgical footage that covers 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (see more details on these surgical techniques in the supplementary material).
3.1 Dataset Curation
We detail how we collected, curated, and annotated these videos. An overview of the pipeline is shown in Fig. 2.
Video collection. We first compiled a list of 35 distinct minimally invasive surgical procedure types [39] (see Fig. 1 in supplementary material). Inspired by works in other fields [14, 9, 56, 1, 41], we collected our videos from YouTube. For each of the 353535 procedures, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively. We run the video collection process until we had 500500500 different videos collected for each procedure leading to a total of $\approx$18K raw videos.
Video classification. The raw videos collected based on search keywords contained a diverse range of content, including patient testimonials and conference presentations. To avoid including non-surgical videos in our dataset, we classified the videos as surgical and non-surgical (Fig.  2b). First, we created 4$\times$4-image video storyboards (i.e., a single image containing 16 key video frames) for each video using the method described in [16]. Second, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig. 2b) was annotated. Third, a binary video storyboard classifier (i.e., a ResNet18 [20]) was trained on this dataset and run over all the raw videos. Finally, manual verification was performed on the videos classified as surgical.
Video selection and trimming. A video classified as surgical may still contain non-surgical frames (e.g., introductory and conclusion slides). Experimentally, we found that the start and end of the surgical footage can be reliably identified by finding the first and last three consecutive surgical frames (sampling the video at 1 fps). Thus, we manually annotated a dataset of 7967 frames (5481 surgical and 2486 non-surgical), trained a surgical frame classifier (i.e., a ResNet18) on it, and used this model to: 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig. 2c).
Video preprocessing. As the videos were from YouTube, we cannot assume that all frames within the trimmed videos are entirely free from non-surgical content (e.g., intraoperative out-of-body views, medical annotations, user-interface details, equipment logos). Therefore, we used our trained frame classifier to detect and obliterate the intraoperative non-surgical frames (Fig. 2d). Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 model [31] to detect and obliterate the non-surgical content in surgical frames. After this process, we manually quality controlled all the remaining curated videos.
Video annotation. We cross-referenced the video titles with the predefined list of procedures. When exact title matches were not found, we leveraged the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt (see Fig. 2 in supplementary material). All annotations including the procedure type and the surgery type (robotic or non-robotic) were manually reviewed and validated to ensure the accuracy of the labels assigned.
Table 2: Linear probing results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks. The experiments are conducted with freezing the backbone and only fine-tuning a linear classification head to demonstrate the foundation model’s effectiveness. The predictions are computed on a frame-by-frame basis for all tasks.
Method Pretraining data Phase recognition Tool detection Action recognition
AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
Acc/F1 Acc/F1 Acc/F1 mAP mAP mAP
MAE [21] LEMON 35.5/32.035.5/32.035.5 / 32.0 54.9/43.454.9/43.454.9 / 43.4 40.7/38.140.7/38.140.7 / 38.1 30.630.630.6 45.245.245.2 24.824.824.8
Endo-FM [53] Private data 51.5/43.151.5/43.151.5 / 43.1 62.7/53.962.7/53.962.7 / 53.9 53.8/49.853.8/49.853.8 / 49.8 50.050.050.0 64.364.364.3 26.226.226.2
EndoViT [5] Merged public data 45.4/38.745.4/38.745.4 / 38.7 46.9/39.046.9/39.046.9 / 39.0 38.5/37.238.5/37.238.5 / 37.2 34.734.734.7 47.147.147.1 29.529.529.5
GSViT [48] GenSurgery 22.0/17.522.0/17.522.0 / 17.5 20.2/13.520.2/13.520.2 / 13.5 22.4/18.022.4/18.022.4 / 18.0 18.218.218.2 35.635.635.6 16.516.516.5
SurgeNetXL [25] SurgeNetXL 68.8/57.068.8/57.068.8 / 57.0 73.2/65.173.2/65.173.2 / 65.1 67.8/61.967.8/61.967.8 / 61.9 72.172.172.1 62.762.762.7 45.345.345.3
LemonFM LEMON $\mathbf{76.4/66.9}$ $\mathbf{75.8/68.6}$ $\mathbf{68.4/62.4}$ $\mathbf{76.8}$ $\mathbf{76.4}$ $\mathbf{50.4}$
Figure 4: Proposed augmented distillation method. To encourage invariance in LemonFM to minor surgical motion and subtle appearance changes across patients, we introduce $W_{i}$ (Eq. 1). The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b). To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video. When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
3.2 Ground truth and proposed new downstream tasks
Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options; and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our LemonFM foundation model, video classification/retrieval, video similarity measurement, and procedure-specific estimation of surgical remaining time).
4 Proposed LemonFM Foundation Model
In this section, we describe how we build LemonFM, our foundation model pretrained on LEMON using a novel augmented distillation approach. Following the standard practice in the field of surgical computer vision [51, 52, 43], we extract frames from the 4194 videos in LEMON at 1 fps for our foundation model pretraining.
Proposed augmented distillation method. A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations). This is to encourage the model to produce similar embeddings for frames containing the same surgical tools and exhibiting comparable tissue appearances.
Inspired by self-supervised knowledge distillation [8], we introduce a novel augmented distillation approach that leverages LEMON and employs ConvNeXt-L [36] for both the teacher and student networks. Specifically, we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$. The student’s parameters $\boldsymbol{\theta_{s}}$ are updated by gradient descent, while the teacher’s parameters $\boldsymbol{\theta_{t}}$ are updated via exponential moving average from the student. We use the outputs of both networks to build two categorical distributions, $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ and $P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))$, where $\sigma_{z}$ denotes the $z^{th}$ element of the softmax output, and $z\in\{1,\dots,C\}$. In training, we minimize the negative log-likelihood of $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))$ following
$\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
where $d(\boldsymbol{u},\boldsymbol{v})=\log\sum_{k^{\prime}=0}^{C-1}\exp\left(\sigma_{k^{\prime}}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))\right)$, $n(\boldsymbol{u},\boldsymbol{v})=\sigma_{\arg\max_{z}P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{u})))}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))$, $I$ is the number of images in the dataset, $U_{i}$ contains two global augmentations of $\boldsymbol{x_{i}}$, and $V_{i}$ contains two global augmentations and four augmented crops of $\boldsymbol{x_{i}}$. Additionally, we introduce $W_{i}$ as an extra supervisory signal which contains two images, and four augmented crops of these two images. The procedure for selecting these two images is depicted in Fig. 4. We use the $k$-nearest neighbors (in embedding space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the adjacent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
Table 3: Full fine-tuning results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks. All experiments are conducted in a full fine-tuning setting (i.e., both backbone and head are fine-tuned). Type S denotes a specialist model trained for a specific task, whereas F denotes a foundation model whose generalization is assessed across multiple surgical downstream tasks. The superscript indicates the use of the standard 10-second relaxed boundary in phase transitions [51]. We also report the performance of LemonFM when trained on only 50% of the labeled training data for the downstream tasks.
Method Type Phase recognition Tool detection Action recognition
AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
Acc/Jacc Acc/Jacc Acc/Jacc mAP mAP mAP
TeCNO [12] S 77.3/50.777.3/50.777.3 / 50.7 88.6/75.188.6/75.188.6 / 75.1 86.1/74.486.1/74.486.1 / 74.4 -- -- --
Trans-SVNet [30] 78.3/50.778.3/50.778.3 / 50.7 90.3/79.390.3/79.390.3 / 79.3 87.2/74.787.2/74.787.2 / 74.7 -- -- --
LoViT [35] 81.4/55.981.4/55.981.4 / 55.9 92.4/81.292.4/81.292.4 / 81.2 /-/-- / - -- -- --
MAE [21] F 63.5/42.263.5/42.263.5 / 42.2 84.6/71.284.6/71.284.6 / 71.2 77.8/63.677.8/63.677.8 / 63.6 72.372.372.3 53.053.053.0 33.133.133.1
Endo-FM [53] 80.3/55.380.3/55.380.3 / 55.3 86.9/76.786.9/76.786.9 / 76.7 79.1/69.979.1/69.979.1 / 69.9 88.488.488.4 84.284.284.2 54.754.754.7
EndoViT [5] 76.1/54.276.1/54.276.1 / 54.2 86.7/75.386.7/75.386.7 / 75.3 78.5/70.078.5/70.078.5 / 70.0 83.983.983.9 77.477.477.4 52.252.252.2
GSViT [48] 60.5/37.060.5/37.060.5 / 37.0 75.3/57.875.3/57.875.3 / 57.8 55.8/39.655.8/39.655.8 / 39.6 34.334.334.3 37.737.737.7 26.326.326.3
SurgeNetXL [25] 80.0/55.180.0/55.180.0 / 55.1 84.4/72.084.4/72.084.4 / 72.0 68.5/58.868.5/58.868.5 / 58.8 86.586.586.5 83.883.883.8 57.557.557.5
LemonFM (ours) - 50% shot 84.9/62.484.9/62.484.9 / 62.4 92.1/78.792.1/78.792.1 / 78.7 86.1/77.786.1/77.786.1 / 77.7 91.591.591.5 89.489.489.4 60.160.160.1
LemonFM (ours) $\mathbf{85.5/64.8}$ $\mathbf{92.7/85.1}$ $\mathbf{89.9/79.4}$ $\mathbf{93.7}$ $\mathbf{94.4}$ $\mathbf{61.9}$
Proposed video classification model. Considering that each surgical procedure is localized to a specific region of the body, we posit that a characteristic scene can be associated with each procedure. Based on this assumption, and inspired by active learning [19], we propose LemonFM-Vid. This is a video classification head that employs the frame embeddings produced by LemonFM to solve our proposed video classification downstream tasks. It does so by aggregating each frame’s embedding following
$\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
where $\boldsymbol{v_{e}}$ is a video embedding, $J$ is the number of frames in the video, $\omega_{j}=\frac{t(\boldsymbol{\phi}_{j})}{\epsilon+\sum_{m=0}^{J-1}t(\boldsymbol{\phi}_{m})}$, $\boldsymbol{\phi}_{j}$ is the embedding of frame $j$, $t(\boldsymbol{\phi})=\left(\frac{1}{K}\sum_{\boldsymbol{\eta}\in K-NN(\boldsymbol{\phi})}d(\boldsymbol{\phi},\boldsymbol{\eta})\right)^{-1}$ is the typicality [19] of a frame, $K-NN(\boldsymbol{\phi})$ with $K=20$ are the nearest neighbors of $\boldsymbol{\phi}$, $d$ is the cosine distance, and $\epsilon=10^{-8}$. Then, each $\boldsymbol{v_{e}}$ is classified with a single-layer MLP.
5 Experiments
5.1 Downstream Tasks, Datasets and Evaluation
We evaluate LemonFM on four surgical downstream tasks: surgical phase recognition [52, 51, 50], surgical tool presence detection [51, 4], surgical action recognition [43], and surgical semantic segmentation [23]. To do so, we use the most widely adopted datasets, data splits, and evaluation metrics.
Surgical phase recognition. For this task, we use three datasets: AutoLaparo [52], M2CAI16 [51], and Cholec80 [51]. These datasets include complete surgical workflows with frame-level phase annotations, where each frame corresponds to a single phase. We perform experiments with both linear probing, where we freeze the backbone and only fine-tune a linear classification head, and full fine-tuning, where both the backbone and the head are fine-tuned. We report the frame-wise accuracy and F1-score for the linear probing setting (Table 2), and the video-level accuracy and Jaccard (more details in supplementary material) for the full fine-tuning setting. To produce video-level online predictions, we replace the linear head with an MS-TCN head [12] (Table 3).
Surgical tool presence detection. Cholec80 and GraSP [4] are utilized in the surgical tool presence detection task. These datasets contain surgical tool presence labels for each frame, allowing multiple tools to be present in a single frame. We report mean Average Precision (mAP) across all frames for both the linear probing and full fine-tuning settings (see Table 2 and 3). We use a linear head for both experiments.
Surgical action recognition. We use CholecT50 [43] as the downstream dataset. Surgical action labels are annotated at frame level, enabling multiple actions occur simultaneously. We use a linear head for both the linear probing and full fine-tuning settings and apply the same metrics as in the surgical tool presence detection task.
Surgical semantic segmentation. We employ the CholecSeg8k [23] dataset, which contains pixel-wise semantic annotations for each frame. The experiments are conducted in full fine-tuning using a feature pyramid network decoder [32]. Dice score is reported over all frames in the dataset (see Table 4).
Table 4: Evaluation on surgical semantic segmentation. The experiments are conducted with using a feature pyramid network (FPN) decoder [32] in a full fine-tuning manner. The superscript denotes the LemonFM backbone is frozen. We also report the performance of LemonFM when trained on only 50% of the labeled training data.
Method Pretraining data CholecSeg8k
mDice
MAE LEMON 68.568.568.5
Endo-FM Private data 54.554.554.5
EndoViT Merged public data 73.073.073.0
GSViT GenSurgery 53.453.453.4
SurgeNetXL SurgeNetXL 70.070.070.0
LemonFM LEMON 71.971.971.9
LemonFM - 50% shot LEMON 78.278.278.2
LemonFM LEMON $\mathbf{81.3}$
5.2 Comparison with SotA
The performance of our LemonFM compared to the state of the art in each downstream task and dataset is shown in Fig. 1, and Tables 2, 3 and 4. These results demonstrate that the diversity of procedures in LEMON enables LemonFM to effectively extract features that are useful for different surgical procedures and downstream tasks.
Surgical phase recognition. As shown in Table 3, LemonFM outperformed every other surigcal foundation model in AutoLaparo ($>$9pp Jacc.), Cholec80 ($>$8pp Jacc.), and M2CAI16 ($>$9pp Jacc.). Furthermore, it also surpassed state-of-the-art specialist models for these datasets. The advantage of LemonFM persists under the linear probing setting (Table 2), showing that the pretrained features are inherently informative and transfer well without additional fine‑tuning.
Surgical tool presence detection. As shown in Table 2 and 3, LemonFM outperformed all surgical foundation models in Cholec80 and GraSP.
Surgical action recognition. LemonFM outperformed all surgical foundation models in CholecT50 (Table 2 and 3).
Surgical semantic segmentation. As shown in Table 4, LemonFM surpassed every other foundation model in the pixel-wise semantic segmentation task by a large margin ($>$8pp mDice).
Low data regime. As shown in Tables 3 and 4, when trained on 50% of the data, LemonFM outperforms other surgical foundation models fine-tuned on 100% of the data, demonstrating high data efficiency.
5.3 Ablation Studies
We conducted incremental ablation studies on the surgical phase recognition (AutoLaparo) and surgical semantic segmentation (CholecSeg8k) tasks (see Table 5).
Table 5: Ablation studies. We froze the backbone and only fine-tuned the task-specific heads (i.e., a linear head for phase recognition and a FPN head for semantic segmentation). The column “Curation” marks whether the raw surgical videos were processed with the pipeline in Fig. 2. The column “Aug” shows whether the augmented distillation method in Fig. 4 was applied.
Dataset Curation Aug Backbone AutoLaparo CholecSeg8k
Acc/F1 mDice
IN-1K N/A ConvNeXt-L 63.6/53.063.6/53.063.6 / 53.0 64.464.464.4
LEMON ConvNeXt-L 71.7/61.471.7/61.471.7 / 61.4 67.467.467.4
LEMON ConvNeXt-L 75.3/65.975.3/65.975.3 / 65.9 68.768.768.7
LEMON ViT-L 75.6/66.175.6/66.175.6 / 66.1 61.261.261.2
LEMON ConvNext-B 74.7/64.574.7/64.574.7 / 64.5 68.568.568.5
LEMON ConvNeXt-L $\mathbf{76.4/66.9}$ $\mathbf{71.9}$
The significance of LEMON. We compare the effects of using LEMON for pretraining the surgical foundation model against those obtained with ImageNet-1K [13] (row 1 and row 3 in Table 5). The model trained on LEMON outperformed the one pretrained on ImageNet by a large margin (+12.9pp F1-score on phase recognition, +4.3pp mDice on semantic segmentation). This finding suggests that pretraining on LEMON offers a superior approach for surgery-specific foundation model training compared to pretraining on general-purpose datasets like ImageNet.
The importance of rigorous data curation. We analyze the importance of our data curation pipeline by comparing LemonFM against a model pretrained on an uncurated version of LEMON (i.e., the raw surgical YouTube videos without our data curation). Our data curation pipeline (Fig. 2) yields substantial improvements. Specifically, the curated dataset leads to an increase of 4.5pp F1-score in the phase recognition task and 1.3pp of mDice in the semantic segmentation task (rows 2 and 3 in Table 5).
The effectiveness of the augmented distillation. We compared the effectiveness of employing our proposed augmented distillation method (Fig. 4) against a vanilla DINO [8]. LemonFM consistently outperformed the vanilla DINO baseline, with the largest gain (+3.2pp mDice) in semantic segmentation (rows 3 and 6 in Table 5).
LemonFM design choices. We performed ablation studies to evaluate the effects of two design choices for surgical foundation model pretraining: backbone architecture, ViT vs. ConvNeXt, and self-supervision method, generative-based [21] vs. discriminative-based [8]. As shown in row 4 and 6 of Table 5, when pretrained on LEMON with our augmented knowledge distillation, the ConvNeXt backbone consistently outperformed the ViT-based counterpart. The improvement in performance is particularly notable in the pixel-wise semantic segmentation task (+10.7pp mDice). In Tables 2, 3, and 4, we compare methods whose pretext task is generative-based (MAE and EndoViT) against methods whose pretext task is discriminative-based (classification) such as Endo-FM, SurgeNetXL and our LemonFM. The results show that discriminative‑based models yield a substantial performance advantage, especially when the backbone is frozen, indicating generative objectives such as MAE are less effective at capturing surgical visual cues.
Table 6: Leaderboard for the proposed tasks on LEMON. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
Method Procedure type Surgery type
mAP/F1 Acc/F1
SlowFast 22.0 /23.9 88.5 /87.5
TimeSformer 42.1/37.5 93.2 /92.7
MViTv2 49.5/41.8 95.8 /94.6
Video Swin Transformer 51.4/47.9 98.8/98.7
LemonFM-Vid $\mathbf{57.8/49.3}$ $\mathbf{98.9/98.9}$
5.4 Leaderboard
We benchmarked state-of-the-art video classification models on LEMON (Table 6), including SlowFast [15], TimeSformer [6], MViTv2 [34] and Video Swin Transformer [37]. Leveraging the robust representations extracted from LemonFM, our video classification model, LemonFM-Vid (Sec. 3), improved multi‑label procedure classification by 6.4pp mAP over all leading models and matched the top performance of the Video Swin Transformer on the binary surgery type classification task. The incorporation of frame-level features extracted by LemonFM, in conjunction with the weighted feature embeddings informed by frame typicality, led to substantial performance gains, underscoring the value of LemonFM for surgical video analysis tasks. Nonetheless, the mAP score of 57.8% achieved by LemonFM-Vid (the top performer method) highlights the challenges associated with our newly proposed procedure type video classification task, underscoring the need for continued innovation in this area.
6 Conclusion
Building on a large corpus of publicly available surgical videos, in this work we presented three main contributions: 1) we proposed a novel data curation pipeline to curate online surgical videos; 2) leveraging this pipeline, we created a large-scale and diverse surgical dataset, LEMON; and 3) we developed a foundation model specific for surgical applications, LemonFM. Our results show that LemonFM outperformed both state-of-the-art surgical foundation models and task-specific architectures across diverse surgical downstream tasks. Even when fine-tuned with only half of the labeled data, LemonFM maintained its advantages. These findings demonstrate LemonFM’s effectiveness and verify the quality of LEMON as a valuable resource for computer vision research in surgery. We expect that both LEMON and LemonFM will serve as a foundation for future research in surgical computer vision and facilitate the advancement of autonomous robotic surgery systems. Future work will concentrate on the development of a surgery-specific video foundation model.
[…]
Supplementary Material
Figure 1: Diversity and procedure prevalence in LEMON. Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right). Distribution of surgical frames by procedure type (center).
Appendix A Surgical Background
In contrast to traditional open surgery (not the target of our work), which entails extensive tissue disruption via large incisions and relies on direct visual inspection by the surgeon (i.e., no cameras are used), modern surgical practices predominantly employ minimally invasive techniques (the target of our work). These techniques involve the insertion of slender instruments, including one that is a camera, into the patient’s body through small incisions.
To control the instruments from outside the patient’s body, there are two variants: robotic-assisted surgery (called “robotic” in our manuscript) and non-robotic laparoscopy (called “non-robotic” in our manuscript). Robotic-assisted surgery involves a surgeon sitting on a console and controlling (with joysticks) several robotic arms that hold and steer the instruments inside the patient. In contrast, non-robotic (also known as traditional or conventional) laparoscopy requires the surgeon to directly hold and steer the laparoscopic hand-held instruments (i.e., no robotic arms involved).
Appendix B Data Curation Details
This section provides details of our data curation pipeline and presents LEMON’s procedure diversity and prevalence (Fig. 1).
B.1 Video Classification
Video summarization. To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e., a single image containing a collage of key video frames) for all the collected videos using the method described in [16]. Storyboards enabled us to quickly determine whether a video contained substantial surgical footage, thereby avoiding the need for complex analysis on the entire sequence.
Video storyboard classification. We manually annotated a dataset comprising 2160 surgical and 1910 non-surgical storyboards. The annotation criterion adopted for labeling a storyboard as surgical was that at least 50% of the key frames contained shots from a surgical camera; specifically, open surgery videos were categorized as non-surgical. To classify the rest of the collected videos, we trained a ResNet18 [20]. To ensure the accuracy of the inference results, we manually reviewed all the videos classified as surgical.
[…]
Table 1: Performance of the video storyboard classification models across five folds.
Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
Fold 0 94.42 94.50 95.88 93.00
Fold 1 93.60 93.50 92.23 95.00
Fold 2 93.89 93.75 91.87 96.00
Fold 3 96.50 96.50 96.50 96.50
Fold 4 95.36 95.50 98.40 92.50
Average 94.75 94.75 94.98 94.60
Std Dev 1.10 1.16 2.66 1.57
B.2 Video Selection and Trimming
Frame classification. A ResNet18 [20] was trained for the surgical/non-surgical video frame classification task. To produce the annotations, the videos were sampled at one frame per second (fps). We annotated 7967 frames, 5481 of which turned to be surgical, and 2486, non-surgical.
Video trimming. Most online videos contain introductory and conclusion slides. Experimentally, we found that the start and end of the surgical footage can be reliably identified by finding the first and last three consecutive frames classified as surgical by our surgical frame classifier (sampling the video at 1 fps). Therefore, this is the approach we used to discard the non-surgical parts at the beginning and end of the collected videos. The resulting videos were manually quality controlled.
[…]
B.3 Video Preprocessing
We use our trained frame classifier to detect and remove the intraoperative non-surgical frames. Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and obliterate the non-surgical content in surgical frames. After this process, we manually quality control all the remaining curated videos.
[…]
You are a highly knowledgeable assistant specializing in surgical procedures and medical terminology.
Your expertise includes identifying and categorizing surgical interventions based on clinical descriptions and procedural contexts.
Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy, splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy, abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass, duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy, appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication, adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal anastomosis, kidney transplant, vaginectomy.
Based on the description of the surgical video: <video title>, determine the most likely procedure type from the list. Focus on matching the description to the procedure type that best aligns with the terminology and context provided.
Figure 2: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
B.4 Video Annotation
For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci. The search for these terms in the video titles was case-insensitive. The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures. For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig. 2. All annotations are manually quality controlled.
[…]
Table 2: Performance of frame classification models across five folds.
Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
Fold 0 94.65 94.85 98.37 91.21
Fold 1 96.43 96.48 97.93 94.98
Fold 2 96.53 96.61 98.69 94.47
Fold 3 94.44 94.60 97.33 91.71
Fold 4 96.15 96.23 98.17 94.22
Average 95.64 95.75 98.10 93.32
Std Dev 0.94 0.88 0.49 1.39
Appendix C LemonFM Pretraining Details
This section provides an overview of the pretraining process for the LemonFM model.
LemonFM pretraining. For LemonFM pretraining, we ran on an Ubuntu 22.04.5 LTS node with eight NVIDIA V100 GPUs (32GB each), using a batch size of 24 per GPU under PyTorch 2.5.1+cu124 (CUDA 12.4). We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), a minimum learning rate of 1e-6, and a random seed of 30. The model was trained for 60 epochs, including 10 warm-up epochs. This number of epochs enables the model to converge and stabilize its training loss on our dataset and pretext task. The model with the lowest training loss was selected as the final model.
Appendix D Downstream Task Details
In this section, we provide details on the datasets used for evaluating each downstream task, including their respective data splits, as well as training configurations for linear probing and full fine-tuning settings. The experiments were conducted on an Ubuntu 22.04.5 LTS node equipped with an Intel IceLake Xeon CPU (72 vCPUs) and two NVIDIA A100 GPUs (each with 80 GB of memory), using PyTorch 2.5.1+cu124 (CUDA 12.4).
D.1 Data splits
Surgical phase recognition. For AutoLaparo, we followed the standard split of ten training videos, four for validation, and seven for testing [52]. For M2CAI16, we followed the data splits in [26, 51, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively. For Cholec80, we adopted the data splits in [51, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing.
Surgical tool presence detection. For GraSP, we adopted the data splits specified in [4], with four videos for training, four for validation, and five for testing. The data split for Cholec80 is identical to those employed in the surgical phase recognition task.
Surgical action recognition. For CholecT50, we followed the formal data split as proposed in [43].
Surgical semantic segmentation. For CholecSeg8k, we followed previous works [25, 18] and use 75% of the videos for training and 25% of the videos for testing (videos 12, 20, 48 and 55).
D.2 Linear probing
We used the teacher model from LemonFM as our backbone. The images were resized to $\times 224$ 224 for all downstream tasks. The downstream models with linear heads were trained with a batch size of 512, an initial learning rate (LR) of 1e-3, the AdamW optimizer, random seed of 30, and cross-entropy loss. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
Table 3: Performance of bounding box detection models across five folds.
Fold Images Instances Precision Recall mAP50 mAP50-95
Fold 0 272 447 74.79 77.63 75.50 62.00
Fold 1 272 438 70.73 77.40 76.36 61.79
Fold 2 272 509 83.34 79.60 84.89 68.56
Fold 3 272 422 75.27 83.69 80.31 68.98
Fold 4 272 496 81.32 81.45 82.58 68.65
Average 77.09 79.94 79.29 66.18
Std Dev 4.41 2.33 3.40 2.98
D.3 Full fine-tuning
We used the teacher model from LemonFM as our backbone and employed three different protocols for downstream task fine-tuning. The images were resized to $\times 224$ 224 for all downstream tasks. The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, random seed of 30, and cross-entropy loss. We use fivefold cross-validation for 50% shot fine-tuning. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
Surgical phase recognition To compare with other specialist models that are specifically tailored for this task, we adapt the video-level accuracy and Jaccard in the full fine-tuning setting. The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames. The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27]. The Jaccard is computed for each class and video, then averaged: first within videos, and secondly across classes [28, 29, 35, 27].
During training, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12]. In the first stage, we trained the backbone with a linear head to perform frame-wise classification without temporal context. Then, we employed the TCN head to incorporate temporal information of the extracted features for predictions. The default TCN configuration was used, consisting of two stages, each containing nine layers. The TCN head was trained using the Adam optimizer with an initial learning rate of 5e-3 for 70 epochs to ensure sufficient training. The checkpoint with the lowest validation loss was selected as the optimal checkpoint.
Appendix E Evaluation of LemonFM-Vid
To evaluate the performance of our video classifier, LemonFM-Vid, we have split the LEMON dataset by videos, with 3369 videos for training and 825 videos for testing. Our vanilla LemonFM was trained on all the LEMON frames, therefore, to evaluate the video classification performance fairly, we have trained a new LemonFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison.
Appendix F Dataset and Licensing
We adhere to the same practices as other datasets created from YouTube and various sources on the Web, such as ImageNet [13], Kinetics [9], YouTube-VIS [56], YouTube-8M [1], Insect-1M [42], Moments-in-Time [41], Tai-Chi-HD [49], HD-Villa-100M [55], and AVSpeech [14]. Specifically, we provide the code to download and generate the LEMON dataset, a list of links to the original YouTube videos, and the corresponding annotations. Researchers working in academic institutions can request direct access to the LEMON dataset (1 fps version) in LMDB format for non-commercial purposes. We will also provide an online form for the YouTube video authors to opt out of our LEMON dataset. The LEMON dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.